# Transfer Learning of Surrogate Models via Domain Affine Transformation Across Synthetic and Real-World Benchmarks

### v1

@@ V-B3 Real-World Optimization Benchmark in Vehicle Dynamics
Figure 5: The comparison evaluates the performance of Random forest regression models obtained through transfer learning against those trained from scratch on the Real-World Optimization Benchmark in Vehicle Dynamics. This analysis examines various transfer data sample sizes. Each cell in the figure shows the percentage difference in average SMAPE (%) between the two approaches for specific transfer settings and sample sizes. Positive values indicate that the transferred model achieves superior accuracy by yielding a lower SMAPE compared to the model trained from scratch.
@@ VI Conclusion
We introduce a transfer learning method specifically designed for non-differentiable surrogate models like Random forest regression (RFR). This method addresses the challenge of domain shifts between source and target functions connected through an unknown affine transformation. By leveraging a small transfer dataset sampled from the target domain, the method optimizes the transformation to adapt a source-trained surrogate model for effective approximation of the target function. This method is particularly valuable in scenarios where data collection is costly, offering superior sample efficiency compared to training a new model from scratch.
Experiments on synthetic scenarios using BBOB test functions demonstrate the effectiveness of the proposed method. With transfer datasets comprising 50–100 samples, the transferred RFR consistently outperformed models trained from scratch, particularly in high-dimensional settings. However, for highly complex functions such as F16 and F21–F24, the benefits of transfer learning are limited, as the original RFR struggles to approximate these functions accurately due to their inherent complexity.
The transfer learning method is further validated on four real-world applications, highlighting its strengths in low-data scenarios. However, challenges emerge when the transfer dataset is extremely small (e.g., 10 samples), leading to potential overfitting. Additionally, substantial differences between source and target domains constrained the model’s adaptability, even with larger datasets. These results underscore the importance of well-aligned transfer datasets to ensure successful adaptation.
Future enhancements could include integrating active learning to prioritize informative data points for fine-tuning and refining both the surrogate model and the transformation simultaneously. While this study demonstrates the feasibility of leveraging pre-trained RFR models in affine transformation scenarios, future work will explore non-linear transformations, such as input-output warping [42], to extend the method’s applicability to more complex and diverse problems.

### v3

@@ V-B3 Real-World Optimization Benchmark in Vehicle Dynamics
Figure 6: The comparison evaluates Random forest regression models obtained through transfer learning against those trained from scratch on F5 of the MarioGAN Suite. It examines various transfer data sample sizes, with each figure cell showing the percentage difference in average SMAPE (%) between the two approaches for specific transfer settings. Positive values indicate better accuracy from transfer learning (i.e. lower SMAPE).
@@ VI Conclusion
We propose a transfer learning method tailored for non-differentiable surrogate models like random forest regression (RFR). This method addresses the challenge of domain shifts between source and target functions connected through an unknown affine transformation. Using a small transfer dataset from the target domain, the method optimizes this transformation to adapt a source-trained surrogate. It offers improved sample efficiency over training from scratch, making it well-suited for data-scarce scenarios. Experiments on BBOB functions show that the transferred RFR outperforms scratch-trained models with just 50–100 transfer samples, especially in high dimensions. However, for complex functions like F16 and F21–F24, transfer learning is less effective due to poor source model accuracy. We further validate the method on four real-world tasks, demonstrating its strength in low-data settings. However, very small transfer sets (e.g., 10 samples) risk overfitting, and large domain shifts limit adaptability even with more data. These findings highlight the need for well-aligned transfer datasets.
Future work could extend this approach to other non-differentiable surrogate models and integrate active learning to jointly refine the surrogate and transformation using informative samples. While our method performs well under affine shifts, its effectiveness drops in cases with complex, nonlinear domain changes (e.g., Vehicle Dynamics). Exploring nonlinear transformations, such as input-output warping [42], may broaden applicability to more challenging real-world tasks.